OpenStudy (anonymous):

Please help! Simplify the expression. Write the answer using scientific notation. 8(8.8 x 10^12) A. 70.4 x 10^12 B. 70.4 x 10^24 C. 70.4 x 10^13 D. 1.68 x 10^13

OpenStudy (ganpat):

A.. 70.4 x 10 `12... As inside the bracket is also multiplication sign.. no need to multiply to both the terms..

OpenStudy (anonymous):

I put that the first time and it's wrong.. :/

OpenStudy (unklerhaukus):

\[8(8.8 \times 10^{12})\]\[ =(8\times8.8)\times10^{12}\]\[=(8\times8+8\times0.8)\times10^{12}\]\[=(64+6.4)\times10^{12}\]\[=(70.4)\times10^{12}\]\[=7.40\times10^{13}\]

OpenStudy (anonymous):

so c?

OpenStudy (ganpat):

no, still A !!!

OpenStudy (unklerhaukus):

C. 70.4 x 10^13 is not right 7.40 x10^13 is the correct answer

OpenStudy (anonymous):

thank you both!

OpenStudy (unklerhaukus):

*ops 7.04x10^13

OpenStudy (anonymous):

thank u!

OpenStudy (ganpat):

@UnkleRhaukus: that option is not available.. and it is same as A .. right ??

OpenStudy (anonymous):

No it's choice C

OpenStudy (unklerhaukus):

scientific notation only has one numeral before the decimal place

OpenStudy (ganpat):

but option c says C. 70.4 x 10^13 !!